1. Calculate the volume of the original sample needed if you would use a 50 mL volumetric flask, a 100 mL volumetric flask, and a 250 mL volumetric flask to prepare a dilution of 1:25.
1
Expert's answer
2021-12-14T14:58:04-0500

Ratio 1:25 means that there have to be 1+25 = 26 equal parts.

1) The volume of the original sample needed if we would use a 50 mL volumetric flask is 50 / 26 = 1.9 ml.

2) The volume of the original sample needed if we would use a 100 mL volumetric flask is 100 / 26 = 3.9 ml.

3) The volume of the original sample needed if we would use a 250 mL volumetric flask is 250 / 26 = 9.6 ml.